I know I need two equations I am confused on where to begin. Can someone help?
Eighty-six passengers rode in an Amtrack train from Boston to Denver. Tickets for regular coach seats $97. Tickets for sleeper car seats $123. The receipts for the trip totaled $9,304. How many passengers purchased each type of ticket?
Need step by step solution to understand
On Monday, Harold picked up seven soft drinks and five burgers for his office. He paid $10.08. On Tuesday, Melinda picked up eight burgers and six soft drinks for everyone. She paid $12.28. What is the cost of one soft drink? What is the cost of one burger?
Tim Duncan scored 35 points in an NBA basketball game without scoring any 3 point shots. He scored 27 times. He made several free throws worth 1 point each and several regular shots from the floor, which were worth 2 points each. How many free throws did he make? How many 2 point shots did he make?

Answer by Paul(988)

1. x+y=86
y=86-x
97x+123(86-x)=9304
97x+10578-123x=9304
-26x=-1274
(((-26x/26=-1274/-26}}}
x=49, 86-49 = 37 About 49 $97 tickects were purchased and 37 $123 dollar tickets were purchased
--------------------------------------------------------------------------
2. Let the burgers be x and soft drinks be y
7x+5y=10.08
y=2.016-1.4x
8x+6y=12.28
8x+6(2.012.016-1.4x)=12.28
8x+12.096-8.4x=12.28
-0.4x=-0.184
x=0.46
Sub in: y=2.016-1.4(0.46)/// y=1.37
Hence the burger costs $1.37 and the soft crink costs $0.46
---------------------------------------------------------------------------
3.Let the one point shot be x and the 2 point shot be y
x=y=27
y=27-x
x+2(27-x)=35
x+54-2x=35
-x=-19
x=19
y=27-(19)
y=8
Hence he scored 19 one pointers and 8 2 pointers